# Gaussian sequence model: sample complexities of testing, estimation and LFHT

### v1

@@ Abstract
We study the Gaussian sequence model, i.e. $X\sim\mathcal{N}({\boldsymbol{\theta}},I_{\infty})$, where ${\boldsymbol{\theta}}\in\Gamma\subset\ell_{2}$ is assumed to be convex and compact. We show that goodness-of-fit testing sample complexity is lower bounded by the square-root of the estimation complexity, whenever $\Gamma$ is orthosymmetric. We show that the lower bound is tight when $\Gamma$ is also quadratically convex, thus significantly extending validity of the testing-estimation relationship from [GP24]. Using similar methods, we also completely characterize likelihood-free hypothesis testing (LFHT) complexity for $\ell_{p}$-bodies, discovering new types of tradeoff between the numbers of simulation and observation samples.
@@ 1 Introduction
which clearly imply (1). While this intuition supports the general validity of (1), unfortunately neither (2) nor (1) hold generally. Indeed, one can easily find counter-example in [Bar02], see (4) below. Nevertheless, in this work we show that both the $\lesssim$ and the $\gtrsim$ of (1) can hold under rather general assumptions. In particular, both of them hold for all $\ell_{p}$-bodies with $p\geq 2$.
@@ 1.1 Our Contribution
where $D(\Gamma,\varepsilon)$ is the Kolmogorov dimension of $\Gamma$ (see also Prop. 6). Hence, under rather general conditions half of the relationship (1) holds up to polylog factors.
@@
which is an example of a more general class of $\ell_{p}$-bodies [Bar02] with $p=1$. For this specific class, we find out in Proposition 2 and Proposition 3 that
@@
Given this counter-example, can we find a general class of convex sets for which validity of (1) can be guaranteed? It turns out that the answer is positive (upto polylogarithmic factors, at least). This is the content of Theorem 3.2, which shows that if in addition to orthosymmetry one requires quadratic convexity [DLM90], we obtain
$\frac{n_{\mathrm{est}}(\Gamma,\varepsilon)}{\varepsilon^{2}}\gtrsim n_{\mathrm{gof}}^{2}(\Gamma,\varepsilon)\,.$
Therefore, together with (3) for such sets we indeed establish (1). In particular, this latter holds for all $\ell_{p}$-bodies with $p\geq 2$.
Our final set of results is in establishing minimax rates (regions) for the LFHT problem. Recall that in [GP24], it was shown that the optimal testing region of LFHT for Gaussian sequence model over an ellipsoid $\Gamma$ satisfies
@@
In particular, this shows that the tradeoff (5) does not extend to $\ell_{p}$-bodies with $p<2$.
@@ 1.2 Related Works
and they show that the minimax sample complexity of goodness-of-fit testing is the smallest $n$ such that $\rho(n)\leq\varepsilon$. More recently, [WW20] refined sample complexity to make it depend on a specific (rather than worst-case) choice of the mean in the null-hypothesis.
@@ 2 Preliminaries and notation
Suppose orthosymmetric, convex, compact set $\Gamma$ satisfies $D(\Gamma,\varepsilon)\lesssim\varepsilon^{-p}$ for some $10<p<$ 1. Then the coordinate-wise Kolmogorov dimension satisfies
@@ 3.3 Testing-Estimation Equivalence for Quadratically Convex Set
Theorem 3.2.
Suppose $\Gamma$ is an orthosymmetric, compact, convex and quadratically convex set, then we have
$n_{\mathrm{gof}}(\Gamma,6\varepsilon)^{2}\leq\frac{n_{\mathrm{est}}(\Gamma,\varepsilon)}{\varepsilon^{2}}.$
The proof of Theorem 3.2 is deferred to Section A.4.
@@ Acknowledgement
Authors would like to thank Oleg Lepski for many discussions and comments.
@@ A.1 Proofs of Proposition 2 and Proposition 3
Since set $\{(\theta_{1},\cdots,\theta_{D}):\sum_{j=1}^{D}j\cdot|\theta_{j}|\leq 1\}$ is compact, there must exists some $(\theta_{1},\cdots,\theta_{D})$ which maximizes (29). Without loss of generality we assume $|\theta_{j}|\leq\sqrt{1/n+\lambda^{2}}$ (otherwise truncate the value of all $\theta_{j}$ to interval $[-\sqrt{1/n+\lambda^{2}},\sqrt{1/n+\lambda^{2}}]$ and it results in another maximizer). Further if there exists $j_{1}\neq j_{2}$ which both satisfy $0<|\theta_{j}|<\sqrt{1/n+\lambda^{2}}$, then by slightly modifying $\theta_{j_{1}}$ and $\theta_{j_{2}}$ we can make (29) larger, while keeping the parameter still in the set. This contradicts to the assumption that $(\theta_{1},\cdots,\theta_{D})$ is a maximizer. Therefore, there exists at most one of $1\leq j\leq D$ which satisfies $0<|\theta_{j}|<\sqrt{1/n+\lambda^{2}}$. We let integer $N$ to be the smallest integer such that
@@
$\prod_{i=1}^{d}\left(1+\frac{h^{2}}{4}\cdot(\exp(nr^{2})-1)+\frac{h^{2}}{4}\cdot(\exp(-nr^{2})-1)\right)$
@@
$\sup_{{\boldsymbol{\theta}}\in\Gamma}\sum_{i=1}^{D}\min\left\{\theta_{i}^{2},\frac{2\log\left(\nicefrac{{2D}}{{(n\varepsilon^{2})}}\right)\vee 0+1}{n}\right\}$
@@
$\chi^{2}\left(\mathbb{E}_{P\sim\mu(\nu)}[P^{\otimes n}]\|P_{0}^{\otimes n}\right)\stackrel{{\scriptstyle(i)}}{{\leq}}\prod_{i=1}^{D}\exp\left(\frac{1}{2}n^{2}(\theta_{i}^{*})^{4}\right)-1=\exp\left(\frac{n^{2}}{2}\sum_{i=1}^{D}(\theta_{i}^{*})^{4}\right)-1,$
@@
$\sup_{{\boldsymbol{\theta}}\in\Gamma}\sum_{i\in\mathbb{Z}_{+}\backslash A}(\theta_{i})^{2}\leq 2\varepsilon^{2}.$
@@ A.3 Discussion between Kolmogorov Dimension and Coordinate-wise Kolmogorov Dimension
Suppose set $\Lambda=\{\mathbf{e}_{1},\cdots,\mathbf{e}_{d}\}$ consists of orthogonal unit vectors. Then for any $(d-1)$-dimensional projection $\Pi$, we have
@@
Suppose orthosymmetric, convex, compact set $\Gamma$ satisfies $D(\Gamma,\varepsilon)\lesssim\varepsilon^{-p}$ for some $10<p<$ 1. Then the coordinate-wise Kolmogorov dimension satisfies
@@
A.4 Proof of Theorem 3.2
We present the proof of Theorem 3.2 in this section.
Proof of Theorem 3.2.
We let $D(\Gamma,3\varepsilon)$ to be the Kolmogorov dimension of set $\Gamma$ at scale $3\varepsilon$ (the definition of Kolmogorov dimension is given in Definition 1). We use $\mathcal{S}^{\star}$ to denote the $D(\Gamma,3\varepsilon)$-dimensional subspace which achieves the minimizer in the definition of Kolmogorov dimension Definition 1, and $\Pi_{\mathcal{S}^{\star}}$ is the projection into $\mathcal{S}^{\star}$. According to [Bar02], as long as $n\geq\sqrt{D(\Gamma,3\varepsilon)}/(5\varepsilon^{2})$, there exists a constant $c$ such that the testing scheme
@@
$\mathbb{P}_{X\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mathbf{0},I)}(\psi(X)=1)\leq\frac{1}{4}\qquad\text{and}\quad\mathbb{P}_{X\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}({\boldsymbol{\theta}},I)}(\psi(X)=0)\leq\frac{1}{4}\quad\text{for any }\|{\boldsymbol{\theta}}\|_{2}\geq 6\varepsilon.$
@@
$n_{\mathrm{gof}}(\Gamma,6\varepsilon)\leq\frac{\sqrt{D(\Gamma,3\varepsilon)}}{\varepsilon^{2}}.$
Next, we notice that according to [DLM90, Page 1428], we have the following lower bound on the density estimation error with $n$ samples:
$\inf_{\widehat{\boldsymbol{\theta}}}\sup_{{\boldsymbol{\theta}}\in\Gamma}\mathbb{E}\left[\left\|\widehat{\boldsymbol{\theta}}(X)-{\boldsymbol{\theta}}\right\|_{2}^{2}\right]$ $\geq\frac{1}{4.44}\cdot\inf_{\Pi}\sup_{{\boldsymbol{\theta}}\in\Gamma}\mathbb{E}\left[\left\|\Pi(X)-{\boldsymbol{\theta}}\right\|_{2}^{2}\right]$
$\geq\frac{1}{4.44}\cdot\inf_{d}\left\{\inf_{\Pi_{d}}\sup_{\theta\in\Gamma}\left\{\|\Pi_{d}(\theta)-\theta\|_{2}^{2}+\frac{d}{n}\right\}\right\}.$
where $\inf_{\Pi}$ denotes infimum over all projections. Hence when $n\leq D(\Gamma,3\varepsilon)/(5\varepsilon^{2})$, for any integer $d<D(\Gamma,3\varepsilon)$,
$\inf_{\Pi_{d}}\sup_{\theta\in\Gamma}\left\{\|\Pi_{d}(\theta)-\theta\|_{2}^{2}+\frac{d}{n}\right\}\geq\inf_{\Pi_{d}}\sup_{\theta\in\Gamma}\|\Pi_{d}(\theta)-\theta\|_{2}^{2}\geq(3\varepsilon)^{2}>5\varepsilon^{2},$
and when $d\geq D(\Gamma,3\varepsilon)$,
$\inf_{\Pi_{d}}\sup_{\theta\in\Gamma}\left\{\|\Pi_{d}(\theta)-\theta\|_{2}^{2}+\frac{d}{n}\right\}\geq\frac{d}{n}\geq 5\varepsilon^{2}.$
@@
$\inf_{\Pi_{d}}\sup_{\theta\in\Gamma}\left\{\|\Pi_{d}(\theta)-\theta\|_{2}^{2}+\frac{d}{n}\right\}\geq\frac{5\varepsilon^{2}}{4.44}>\varepsilon^{2},$
@@
$n_{\mathrm{est}}(\Gamma,\varepsilon)\geq\frac{D(\Gamma,3\varepsilon)}{5\varepsilon^{2}}\geq\varepsilon^{2}\cdot n_{\mathrm{gof}}(\Gamma,6\varepsilon)^{2}.$

### v4

# Testing and estimation in orthosymmetric Gaussian sequence model

@@ Abstract
We study the Gaussian sequence model, i.e. $X\sim\mathcal{N}({\boldsymbol{\theta}},I_{\infty})$, where ${\boldsymbol{\theta}}\in\Gamma\subset\ell_{2}$ is assumed to be convex and compact. We show that goodness-of-fit testing sample complexity is lower bounded by the square-root of the estimation complexity, whenever $\Gamma$ is orthosymmetric. This lower bound is tight when $\Gamma$ is also quadratically convex (as shown by [DLM90, Ney23]). We also completely characterize likelihood-free hypothesis testing (LFHT) complexity for $\ell_{p}$-bodies, discovering new types of tradeoff between the numbers of simulation and observation samples, compared to the case of ellipsoids ($p=2$) studied in [GP24].
@@ 1 Introduction
which clearly imply (1). In addition to examples from [GP24], the relationship (1) also holds for a special class of Gaussian sequence models with a QCO (compact , convex, quadratically convex and orthosymmetric) constraint set, as follows from [DLM90] and [Ney23] (see Prop. 4 below).
While these considerations support the validity of (1), unfortunately neither (2) nor (1) hold in general. Indeed, a counter-example is implicitly contained in [Bar02], see (4) below. Nevertheless, in this work we show that the $\lesssim$ direction of (1) holds under the mere assumption of orthosymmetry, see (3).
@@ 1.1 Our Contribution
where $D(\Gamma,\varepsilon)$ is the Kolmogorov dimension of $\Gamma$ (see also Prop. 7). Hence, under rather general conditions half of the relationship (1) holds up to polylog factors. The unusual aspect of our proof is that a lower bound on testing is shown by extracting a hard to test mixture distribution from the analysis of a soft-thresholding estimator.
As we mentioned above, if in addition to orthosymmetry one also assumes quadratic-convexity of $\Gamma$ (i.e. if $\Gamma$ is QCO), then results of [DLM90] on estimation and [Ney23] on testing together imply validity of (1) for such sets (see Prop. 4), thus showing that our lower bound is generally tight. We recall that $\ell_{p}$-bodies [Bar02, DLM90] with $p\geq 2$ are QCO sets.
@@
which is an example of a more general class of an $\ell_{p}$-body with $p=1$. For this specific class, we find out in Proposition 2 and Proposition 3 that
@@
Our second contribution is in establishing minimax rates (regions) for the LFHT problem defined in (10). Recall that in [GP24], it was shown that the optimal testing region of LFHT for Gaussian sequence model over an ellipsoid $\Gamma$ satisfies
@@
In particular, this shows that the “regular LFHT” region, where the boundary is defined in terms of the product $mn$ as in (5), is specific to $\ell_{p}$-bodies with $p\geq 2$, while for $p<2$ the region is rather different.
@@ 1.2 Related Works
and they show that the minimax sample complexity of goodness-of-fit testing is the smallest $n$ such that $\rho(n)\leq\varepsilon$. More recently, [WW20] refined sample complexity to make it depend on a specific (rather than worst-case) choice of the mean in the null-hypothesis. In the special case of QCO sets $\Gamma$ Neykov [Ney23] characterizes (within a constant factor) the goodness-of-fit testing sample complexity in terms ofcritical radius, which in turn is derived from Kolmogorov widths. The case of $\Gamma$ being a $d$-dimensional convex cone was studied in [WWG19], in particular demonstrating that in the case of “ice-cream cones” one can get $n_{\mathrm{gof}}\asymp{1\over\varepsilon^{2}}$ but $n_{\mathrm{est}}\asymp{d\over\varepsilon^{2}}$, due to null-case being at the apex of the cone.
@@ 2 Preliminaries and notation
Suppose orthosymmetric, convex, compact set $\Gamma$ satisfies $D(\Gamma,\varepsilon)\lesssim\varepsilon^{-p}$ for some $0<p<1$. Then the coordinate-wise Kolmogorov dimension satisfies
@@ 3.2 One-side inequality for Orthosymmetric Convex Sets
Remark 3.
We notice that without assuming the orthosymmetric property, Theorem 3.1 can fail. For example,  [WWG19] shows that for the $d$-dimensional circular cone with null-hypothesis at the apex of the cone the goodness-of-fit testing complexity is $n_{\mathrm{gof}}\asymp 1/\varepsilon^{2}$ (independent to the dimension $d$), while the estimation complexity is the usual $n_{\mathrm{est}}\asymp d/\varepsilon^{2}$. Based on this idea, a non-orthosymmetric counter-example to the lower bound in the Theorem can be constructed by taking
$\Gamma=\left\{{\boldsymbol{\theta}}=(\theta_{1},\theta_{2},\cdots):\sum_{i=2}^{\infty}\theta_{i}^{2}\leq\theta_{1}^{2},\theta_{1}\geq 0\right\}\cap\left\{{\boldsymbol{\theta}}=(\theta_{1},\theta_{2},\cdots):\sum_{i=1}^{\infty}i^{2}\cdot\theta_{i}^{2}\leq 1\right\}\,.$
Indeed, it can be shown similarly to [WWG19] that $n_{\mathrm{gof}}\asymp\varepsilon^{-2}$, while $n_{\mathrm{est}}\gtrsim\varepsilon^{-3}$.
@@ 3.3 Testing-Estimation Equivalence for Quadratically Convex Set
Proposition 4 ([Ney23, DLM90]).
There exists universal constants $c,C>0$ such that for any QCO (compact, convex, quadratically convex and orthosymmetric) set $\Gamma$,
$n_{\mathrm{gof}}(\Gamma,c\varepsilon)^{2}\lesssim\frac{n_{\mathrm{est}}(\Gamma,\varepsilon)}{\varepsilon^{2}}\lesssim n_{\mathrm{gof}}(\Gamma,C\varepsilon)^{2},\qquad\forall\varepsilon>0.$
Proof.
Seminal work [DLM90] shows that for QCO sets, $D(\Gamma,c\varepsilon)/\varepsilon^{2}\lesssim n_{\mathrm{est}}(\Gamma,\varepsilon)\lesssim D(\Gamma,C\varepsilon)/\varepsilon^{2}$. In [Ney23] it is shown that $n_{\mathrm{gof}}(\Gamma,c\varepsilon)\lesssim\sqrt{D(\Gamma,\varepsilon)}/\varepsilon^{2}\lesssim n_{\mathrm{gof}}(\Gamma,C\varepsilon)$ for QCO sets. These results together imply the proposition. ∎
In Section A.4 we also include a slightly more general result showing that the second inequality in the Proposition holds under the weaker assumption on $\Gamma$ (namely, only requiring that projection estimators be order-optimal).
@@ Acknowledgement
Authors would like to thank S. Balakrishnan, M. Neykov and Y. Wei for pointing out several important results in the literature. We would also like to thank Oleg Lepski and Martin Wainwright for many discussions and comments.
@@ A.1 Proofs of Proposition 2 and Proposition 3
Since set $\{(\theta_{1},\cdots,\theta_{D}):\sum_{j=1}^{D}j\cdot|\theta_{j}|\leq 1\}$ is compact, there must exists some $(\theta_{1},\cdots,\theta_{D})$ which maximizes (29). Without loss of generality we assume $|\theta_{j}|\leq\sqrt{1/n+\lambda^{2}}$ (otherwise truncate the value of all $\theta_{j}$ to interval $\sqrt{1/n+\lambda^{2}},\sqrt{1/n+\lambda^{2}}]$ and it results in another maximizer). Further if there exists $j_{1}\neq j_{2}$ which both satisfy $0<|\theta_{j}|<\sqrt{1/n+\lambda^{2}}$, then by slightly modifying $\theta_{j_{1}}$ and $\theta_{j_{2}}$ we can make (29) larger, while keeping the parameter still in the set. This contradicts to the assumption that $(\theta_{1},\cdots,\theta_{D})$ is a maximizer. Therefore, there exists at most one of $1\leq j\leq D$ which satisfies $0<|\theta_{j}|<\sqrt{1/n+\lambda^{2}}$. We let integer $N$ to be the smallest integer such that
@@
$\hskip-14.22636pt\prod_{i=1}^{d}\left(1+\frac{h^{2}}{4}\cdot(\exp(nr^{2})-1)+\frac{h^{2}}{4}\cdot(\exp(-nr^{2})-1)\right)$
@@
$\hskip-14.22636pt\sup_{{\boldsymbol{\theta}}\in\Gamma}\sum_{i=1}^{D}\min\left\{\theta_{i}^{2},\frac{2\log\left(\nicefrac{{2D}}{{(n\varepsilon^{2})}}\right)\vee 0+1}{n}\right\}$
@@
$\hskip-28.45274pt\chi^{2}\left(\mathbb{E}_{P\sim\mu(\nu)}[P^{\otimes n}]\|P_{0}^{\otimes n}\right)\stackrel{{\scriptstyle(i)}}{{\leq}}\prod_{i=1}^{D}\exp\left(\frac{1}{2}n^{2}(\theta_{i}^{*})^{4}\right)-1=\exp\left(\frac{n^{2}}{2}\sum_{i=1}^{D}(\theta_{i}^{*})^{4}\right)-1,$
@@
$sup_{{\boldsymbol{\theta}}\in\Gamma}\sum_{i\in\mathbb{Z}_{+}\backslash A}(\theta_{i})^{2}\leq 2\varepsilon^{2}.$
@@ A.3 Discussion between Kolmogorov Dimension and Coordinate-wise Kolmogorov Dimension
Suppose set $\Lambda=\{\mathbf{e}_{1},\cdots,\mathbf{e}_{d}\}$ consists of orthogonal unit vectors. Then for any (d1)(d-1)-dimensional projection $\Pi$, we have
@@
Suppose orthosymmetric, convex, compact set $\Gamma$ satisfies $D(\Gamma,\varepsilon)\lesssim\varepsilon^{-p}$ for some $0<p<1$. Then the coordinate-wise Kolmogorov dimension satisfies
@@
A.4 Alternative proof of Proposition 4
We present a version of the second inequality in Proposition 4 in this section.
Proposition 8.
Suppose $\Gamma$ any set such that the projection estimator is minimax optimal up to constants, i.e. there exists a positive constant $c$ such that
$\inf_{\Pi}\sup_{{\boldsymbol{\theta}}\in\Gamma}\mathbb{E}\left[\left\|\Pi(X)-{\boldsymbol{\theta}}\right\|_{2}^{2}\right]\leq c\cdot\inf_{\widehat{\boldsymbol{\theta}}}\sup_{{\boldsymbol{\theta}}\in\Gamma}\mathbb{E}\left[\left\|\widehat{\boldsymbol{\theta}}(X)-{\boldsymbol{\theta}}\right\|_{2}^{2}\right],$ (49)
where $\Pi$ denotes the class of projection estimators, and $\boldsymbol{\theta}$} denotes any estimators. Then we have
$n_{\mathrm{gof}}(\Gamma,2\sqrt{c}\varepsilon)^{2}\leq\frac{81n_{\mathrm{est}}(\Gamma,\varepsilon)}{\varepsilon^{2}}.$
We remind that [DLM90] famously showed that (49) holds for QCO sets. Thus, the result above strictly generalizes second inequality in Proposition 4.
Proof.
We let $D(\Gamma,\sqrt{c}\varepsilon)$ to be the Kolmogorov dimension of set $\Gamma$ at scale $\sqrt{c}\varepsilon$ (the definition of Kolmogorov dimension is given in Definition 1). We use $\mathcal{S}^{\star}$ to denote the $D(\Gamma,3\varepsilon)$-dimensional subspace which achieves the minimizer in the definition of Kolmogorov dimension Definition 1, and $\Pi_{\mathcal{S}^{\star}}$ is the projection into $\mathcal{S}^{\star}$. According to [Bar02], as long as $n\geq 9\sqrt{D(\Gamma,\sqrt{c}\varepsilon)}/(c\varepsilon^{2})$, there exists a constant $c$ such that the testing scheme
@@ A.4 Alternative proof of Proposition 4
$\mathbb{P}_{X\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}(\mathbf{0},I)}(\psi(X)=1)\leq\frac{1}{4}\qquad\text{and}\quad\mathbb{P}_{X\stackrel{{\scriptstyle\text{iid}}}{{\sim}}\mathcal{N}({\boldsymbol{\theta}},I)}(\psi(X)=0)\leq\frac{1}{4}\quad\text{for any }\|{\boldsymbol{\theta}}\|_{2}\geq 2\sqrt{c}.$
@@
$n_{\mathrm{gof}}(\Gamma,2\sqrt{c}\varepsilon)\leq\frac{9\sqrt{D(\Gamma,\sqrt{c}\varepsilon)}}{c\varepsilon^{2}}.$
Next, according to Eq. (49),
$\inf_{\widehat{\boldsymbol{\theta}}}\sup_{{\boldsymbol{\theta}}\in\Gamma}\mathbb{E}\left[\left\|\widehat{\boldsymbol{\theta}}(X)-{\boldsymbol{\theta}}\right\|_{2}^{2}\right]$ $\geq\frac{1}{c}\cdot\inf_{\Pi}\sup_{{\boldsymbol{\theta}}\in\Gamma}\mathbb{E}\left[\left\|\Pi(X)-{\boldsymbol{\theta}}\right\|_{2}^{2}\right]$
$\geq\frac{1}{c}\cdot\inf_{d}\left\{\inf_{\Pi_{d}}\sup_{{\boldsymbol{\theta}}\in\Gamma}\left\{\|\Pi_{d}({\boldsymbol{\theta}})-{\boldsymbol{\theta}}\|_{2}^{2}+\frac{d}{n}\right\}\right\}.$
where $\inf_{\Pi}$ denotes infimum over all projections. Hence when $n<D(\Gamma,\sqrt{c}\varepsilon)/(c\varepsilon^{2})$, for any integer $d<D(\Gamma,\sqrt{c}\varepsilon)$,
$\inf_{\Pi_{d}}\sup_{{\boldsymbol{\theta}}\in\Gamma}\left\{\|\Pi_{d}({\boldsymbol{\theta}})-{\boldsymbol{\theta}}\|_{2}^{2}+\frac{d}{n}\right\}>\inf_{\Pi_{d}}\sup_{{\boldsymbol{\theta}}\in\Gamma}\|\Pi_{d}({\boldsymbol{\theta}})-{\boldsymbol{\theta}}\|_{2}^{2}\geq(\sqrt{c}\varepsilon)^{2}=c\varepsilon^{2},$
and when $d\geq D(\Gamma,\sqrt{c}\varepsilon)$,
$\inf_{\Pi_{d}}\sup_{{\boldsymbol{\theta}}\in\Gamma}\left\{\|\Pi_{d}({\boldsymbol{\theta}})-{\boldsymbol{\theta}}\|_{2}^{2}+\frac{d}{n}\right\}\geq\frac{d}{n}>c\varepsilon^{2}.$
@@
$\inf_{\widehat{\boldsymbol{\theta}}}\sup_{{\boldsymbol{\theta}}\in\Gamma}\mathbb{E}\left[\left\|\widehat{\boldsymbol{\theta}}(X)-{\boldsymbol{\theta}}\right\|_{2}^{2}\right]>\frac{c\varepsilon^{2}}{c}=\varepsilon^{2},$
@@
$81n_{\mathrm{est}}(\Gamma,\varepsilon)\geq\frac{81D(\Gamma,\sqrt{c}\varepsilon)}{c\varepsilon^{2}}\geq\varepsilon^{2}\cdot n_{\mathrm{gof}}(\Gamma,2\sqrt{c}\varepsilon)^{2}.$